OpenStudy (anonymous):

Simplify 5x^2-19x-4/x-4 will give medal

OpenStudy (cwrw238):

first factor the numerator 5x^2 - 19x -4

OpenStudy (cwrw238):

it ill look like (5x +/- a)(x +/- b) can you figure out the a and b and correct signs

OpenStudy (cwrw238):

hopefully one factor will be (x - 4) so we can simplify

OpenStudy (cwrw238):

lets suppose it is x - 4 then we have (5x +/- a)( x - 4) now +/- a will be equal to +1 to get the -4 in the original expression but will that give us the -19x?

OpenStudy (cwrw238):

if so then we got it

OpenStudy (anonymous):

5*-4=-20 20*-1=-20 20-1=19 \[\frac{ 5x ^{2} -19x-4}{ x-4 }=\frac{ 5x ^{2} -\left( 20-1 \right)x-4}{ x-4 }=\frac{ 5x ^{2}-20x+1x-4 }{ x-4 }\] calculate the factors of the numerator and cancel the common factor.